OpenStudy (anonymous):

I need help on a pre-calc problem: Verify the identity. cosine of x divided by quantity one plus sine of x plus quantity one plus sine of x divided by cosine of x equals two times secant of x.

OpenStudy (anonymous):

Here's a better view of it (cos x/ 1 + sin x) + (1 + sin x/ cos x) = 2 sec x

zepdrix (zepdrix):

\[\Large\rm \frac{\cos x}{1+\sin x}+\frac{1+\sin x}{\cos x}=2\sec x\]

zepdrix (zepdrix):

I guess what I would do isssss... Multiply the first fraction by the `conjugate` of the `denominator`. That should get us on the right track. Don't mess with the second fraction for now.

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left(\frac{1-\sin x}{1-\sin x}\right)}\frac{\cos x}{1+\sin x}+\frac{1+\sin x}{\cos x}\]

zepdrix (zepdrix):

Remember what happens when you multiply conjugates? :)

OpenStudy (anonymous):

What do you mean specifically?

zepdrix (zepdrix):

(1-sin x) and (1+sin x) are called conjugates because they are `the same` except for the sign between them. When multiplying conjugates, you end up with the difference of squares:\[\Large\rm (a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

Oh, okay. So you'd end up with a denominator of 1+sinx^2?

zepdrix (zepdrix):

\[\Large\rm\frac{(1-\sin x)\cos x}{1-\sin^2 x}+\frac{1+\sin x}{\cos x}\]Ok great! Let's NOT do the multiplication in the top, just leave it like this for now.

zepdrix (zepdrix):

Hmm, anything we can do with 1-sin^2x? Do you remember your Pythagorean Identity involving sine and cosine?

OpenStudy (anonymous):

sin^2x+cos^2x=1

zepdrix (zepdrix):

Hmmm, yah. Can we use that maybe? If we subtract sin^2x from each side,\[\Large\rm \cos^2x=1-\sin^2x\]Ooo, how bout that?

OpenStudy (anonymous):

Woah, okay this was what I was missing. Okay so next would we dot he same for the second fraction's numerator or?

zepdrix (zepdrix):

No no. Let's leave the second fraction alone. We're about to run into a common denominator almost by accident here.

zepdrix (zepdrix):

\[\Large\rm\frac{(1-\sin x)\cos x}{\cos^2x}+\frac{1+\sin x}{\cos x}\]So we're here, yes? In the first fraction, any common factors between the numerator and denominator that we can cancel out?

OpenStudy (anonymous):

Oh so it'd be \[1-sinx/cosx\]

zepdrix (zepdrix):

\[\Large\rm\frac{(1-\sin x)\cancel{\cos x}}{\cos^{\cancel{2}}x}+\frac{1+\sin x}{\cos x}\] \[\Large\rm\frac{(1-\sin x)}{\cos x}+\frac{1+\sin x}{\cos x}\]Mmm yes, very good!

OpenStudy (anonymous):

Wow, okay, now we have a common denominator so we can add right. and then could we use the sin and cos Pythagorean identity?

zepdrix (zepdrix):

Correct, let's just simply add the tops from there.\[\Large\rm \frac{1-\sin x+1+\sin x}{\cos x}\]No, we won't need to use the Pythagorean Identity again.

zepdrix (zepdrix):

Combine like-terms in the numerator c: What are you left with on top?

OpenStudy (anonymous):

just 2 right?

zepdrix (zepdrix):

\[\Large\rm \frac{2}{\cos x}\]Ok good. Let's write it like this, pulling the 2 in front as multiplication,\[\Large\rm 2\frac{1}{\cos x}\]Another identity should jump out at you ;)

OpenStudy (anonymous):

Awh yes. now we have 2secx. Thank you so so so so much. My algebra skills just fly out the window when we involve identities so thank you.

zepdrix (zepdrix):

Yayyy team \c:/ No problem.